OpenStudy (anonymous):

If I varies inversely with R and I = 0.25 when R = 18, which equation should be used to show this relationship? A.IR = 0.45 B.l=4.5/R C.I = 4.5R D.l=R/4.5

OpenStudy (anonymous):

@carolinagb02

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Inverse relationship means that whatever you have for I and R, the product is ALWAYS THE SAME.

OpenStudy (anonymous):

So then what do I do?

Parth (parthkohli):

The question says that if I = 0.25 then R = 18. That means that IR = 18 * 0.25 = 4.5. But because I and R are inversely related, I*R is ALWAYS 4.5.

Parth (parthkohli):

I * R = 4.5

Parth (parthkohli):

Divide both sides by R

OpenStudy (anonymous):

dont you mean I*R=0.45?

Parth (parthkohli):

What is 18 * 0.25?

OpenStudy (anonymous):

4.5

Parth (parthkohli):

So it is IR = 4.5

OpenStudy (anonymous):

but thats not an option i=r*4.5?

Parth (parthkohli):

Divide both sides by R.

OpenStudy (anonymous):

so l=0.25

Parth (parthkohli):

Uh, nope. I R = 4.5

OpenStudy (anonymous):

ok then what?

Parth (parthkohli):

Divide both sides by R in the equation above

OpenStudy (anonymous):

but how do I do that?

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Divide the left hand side by R and the right hand side by R.

OpenStudy (anonymous):

ok so its l=4.5/R which is 18?

Parth (parthkohli):

Madam, you have already obtained your answer.

OpenStudy (anonymous):

oh omg thanks so much! :)